# APVR: Hour-Level Long Video Understanding with Adaptive Pivot Visual Information Retrieval

### v2

@@ Abstract
Current multimodal large language models (MLLMs) struggle with hour-level video understanding, facing significant challenges not only in modeling the substantial information volume of long videos but also in overcoming the memory wall and resource constraints during both training and inference. Although recent training-free approaches have alleviated resource demands by compressing visual features, their reliance on incomplete visual information limits the performance potential. To address these limitations, we propose Adaptive Pivot Visual information Retrieval (APVR), a training-free framework that hierarchically retrieves and retains sufficient and important visual information. It breakthroughs the memory wall limitation via two complementary components: Pivot Frame Retrieval employs query expansion and iterative spatio-semantic confidence scoring to identify relevant video frames, and Pivot Token Retrieval performs query-aware attention-driven token selection within up to 1024 pivot frames. This dual granularity approach enables the processing of hour-long videos while maintaining semantic fidelity. Experimental validations demonstrate significant performance improvements, achieving 64.9% on LongVideoBench and 68.4% on VideoMME, which are state-of-the-art results for both training-free and training-based approaches. Meanwhile, our method provides plug-and-play integration capability with existing MLLM architectures.
1 Introduction
The proliferation of long-form video content necessitates robust understanding capabilities that extend beyond the current limitations of multimodal large language models (MLLMs). While existing video-based models (Lin et al. 2023; Zhang et al. 2025) demonstrate proficiency on short video sequences, their scalability for hour-level video remains challenging. Specifically, they encounter not only the memory wall but also the inherent dilution of semantic information across extensive temporal spans.
The core challenge stems from the inability of current Video MLLMs to effectively model and extract critical information from long video sequences. In hour-level videos, semantically relevant contents are typically sparse across both temporal and spatial dimensions. However, existing approaches process entire sequences by uniform sampling, leading to computational inefficiency and degraded performance due to attention dispersion across irrelevant segments. As illustrated in Fig. 1(a), this strategy samples frames that are not pivot, leading to wrong answer generated by MLLMs.
@@ 1 Introduction
Recently, some approaches (Ataallah et al. 2024; Tang et al. 2025a; Guo et al. 2025; Ye et al. 2025; Cheng et al. 2025; Luo et al. 2025) have attempted to retrieve query-relevant frames from long videos, simplifying long video understanding task as key-frame selection. Consequently, MLLMs process only few and sparse key frames instead of continuous video frames, neglecting the temporal-semantic relationships and potentially failing in scenarios requiring comprehensive temporal reasoning and long-range logical understanding, as shown in Fig. 1(b).
A potential solution to the above limitation is to retrieve sufficient pivot frames and improve coverage as shown in Fig. 1(c). However, such approaches encounter the memory wall when processing these frames with MLLMs. While recent training-based methods employ strategies such as multimodal sequence parallelism(Fu et al. 2025; Liu et al. 2024d) or feature compression (Shu et al. 2024; Liu et al. 2025a) during the training phase to accommodate extended video frames, they demand resource-intensive multi-stage training based on existing MLLMs. Thus, they lack adaptability to rapidly evolving MLLM architectures due to high retraining costs.
@@
The inherent trade-off between temporal coverage and computational feasibility has hindered existing methods from achieving comprehensive video understanding. To address this core challenge, we propose Adaptive Pivot Visual information Retrieval (APVR) framework, a training-free approach that integrates spatio-semantic frame retrieval with token retrieval and compression, as shown in Fig. 1(d). This two-stage retrieval mechanism enables the system to concentrate computational resources on the pivot visual information that is most relevant to the input query, thus efficiently processing hour-long videos without encountering memory limitations.
@@
Specifically, at the frame level, our Pivot Frame Retrieval (PFR) component expands the original query into four types of semantic information: objects, descriptions, relations, and semantics. This expansion enables more comprehensive frame scoring from spatial and semantic perspectives using complementary visual models: CLIP for semantic similarity and Grounding-DINO for object detection and spatial reasoning. The framework incorporates temporal diffusion mechanisms to maintain temporal coherence and employs adaptive resampling strategies to progressively refine frame selection across multiple iterations. At the token level, our Pivot Token Retrieval (PTR) component extends the concept of semantic importance to fine-grained visual representations. This component leverages query-aware multi-layer attention scoring to identify the most relevant visual tokens within selected frames, employing dynamic chunk-wise selection and head-wise soft voting mechanisms to maintain both computational efficiency and semantic accuracy.
The main contributions of this work are summarized as follows:
@@ 2 Related Work
Video MLLMs  Recent years have witnessed significant advances in multimodal large language models (MLLMs) for visual perception and understanding. As a critical modality in visual data analysis, video has also been integrated into MLLM frameworks. Notable efforts include Qwen2.5-VL (Bai et al. 2025a), InternVideo2.5 (Wang et al. 2025b), VideoLlama3 (Zhang et al. 2025), and VideoChat-Flash (Li et al. 2024b), which explore video understanding through a unified architecture consisting of three key components: (a) a video encoder for spatio-temporal feature extraction, (b) a projector for aligning visual features with linguistic embeddings, and (c) a video-aware LLM for multimodal reasoning. Due to the inherent complexity of video content (e.g., multi-object interactions, long-range temporal dependencies), existing MLLMs often fail to extract semantically salient information for video-centric downstream tasks requiring precise spatiotemporal reasoning, such as video question answering and spatiotemporal video grounding. Some works(Bai et al. 2025a; Wei et al. 2025) exploit to aggregate timestamps in the RoPE(Su et al. 2024). However, these methods exhibit significant performance degradation when extrapolating to videos exceeding their pre-training duration. Moreover, current architectures also exhibit limitations of inefficient cross-modal alignment between visual information and linguistic queries, particularly in long-form videos exceeding 5 minutes (Li et al. 2024a).
@@
Long Video Understanding  Recently advances in long video understanding can be broadly categorized into training-based and training-free paradigms. In the training-based paradigm, MLLMs are directly trained using large-scale long video datasets. The whole training process is usually divided into several stages to stably improve different components of MLLM, as well as to improve the ability to model different types of input visual data (Li et al. 2024b; Zhang et al. 2025). In order to process extra-long videos, current methods explore two different strategies: sequence parallelism (Chen et al. 2024; Zhang et al. 2024b; Shen et al. 2025) or visual feature compression (Song et al. 2024; Shen et al. 2024; Liu et al. 2025a). For instance, LongVILA (Chen et al. 2024) adopts a five-stage training pipeline with a novel multi-modal sequence parallelism to efficiently understand long videos. Video-XL-Pro (Liu et al. 2025a) builds a learnable module to generate compact and comprehensive video tokens for hour-long video understanding. In the training-free paradigm, long video understanding task is organized as an agent-like process (Liu et al. 2025b; Wang et al. 2024c), where MLLMs serve as reasoning engine within the system. These training-free methods mainly focus on retrieving the most important information in video frames (Tang et al. 2025b; Park et al. 2024; Ye et al. 2025; Guo et al. 2025; Cheng et al. 2025) or visual tokens (Wang et al. 2025a; Luo et al. 2025). For example, AKS (Tang et al. 2025b) adopts a plug-and-play module for keyframe selection and information pre-filtering for video-based MLLMs. QuoTA (Luo et al. 2025) adopts an ante-hoc training-free modular for query-aware visual token assignment and important information retrieval. Despite these advancements, few methods simultaneously explore frame and token optimization, which is addressed in this work through adaptive pivot information retrieval.
3 Method
@@
3.1 Overview
The framework of the proposed APVR is illustrated in Fig. 2, which is a training-free system. Given an input video =$\mathcal{F}=\{f_{t}\}_{t=1}^{N}$ with $N$ frames sampled under specific $fps$ and query $Q$, the system aims to retrieve the most important visual information for accurate response. To achieve this, we build two main components in APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTF). In PFR, we search pivot frames $\mathcal{F}_{pivot}$ with high confidence scores using expanded queries and basic visual models. In PTR, we explore an attention-driven token selection in MLLMs to retrieve pivot tokens $\mathcal{T}_{pivot}$ that are most relevant with expanded queries.
3.2 Pivot Frame Retrieval
@@
Objects: The key objects $Obj_{k}$ or cue objects $Obj_{c}$ that are detectable by visual models such as Grounding-DINO (e.g., “bold person”,“phone”,“blue background”’). All objects are aggregated as $Obj=\{Obj_{k},Obj_{c}\}$.
@@
4:$\mathcal{S}\leftarrow\mathbf{0}^{N}$triangleright Confidence scores
5:$\mathcal{V}\leftarrow\emptyset$triangleright Visited frames set
6:$\{Obj,Des,Rel,Sem\}\leftarrow\operatorname{LLM}(Q)$triangleright Query Expansion
@@
$\mathcal{S}_{i}=\mathrm{max}(\mathcal{S}_{i},\frac{\mathcal{S}_{t}}{1+|i-t|}),i\in[t-w,t+w]$ (2)
@@
$\mathcal{H}_{s}=\mathrm{TopK}(\mathcal{S}\odot\mathcal{U},\frac{N}{2\nabla_{p}})$ (3)
@@
$\mathcal{E}_{\gamma}(s_{i})=-\sum_{j=i-\gamma}^{i+\gamma}e_{j}\cdot\text{log}(e_{j}),e_{j}=\frac{s_{j}}{\sum_{k=j-\gamma}^{j+\gamma}s_{k}}$ (4)
, then the frames with high entropy are identified by:
$\mathcal{H}_{e}=\{f_{t}\mid\mathcal{E}_{\gamma}(s_{t})>\mu+0.5\sigma\}$ (5)
, where $\mu$ and $\sigma$ are the mean and standard deviation of the entropy in $\mathcal{C}$.
@@
$\mathcal{C}_{\text{rand}}\sim\mathrm{Unif}(\mathrm{Perm}(S\odot\mathcal{C},\alpha N_{\mathcal{C}}))$ (8)
, where $\alpha$ balances the multinomial score-based importance sampling $\mathcal{C}_{\text{multi}}$ and the diverse random sampling $\mathcal{C}_{\text{rand}}$ to capture the full semantic content of the video and also to avoid falling into local optimal.
@@
Model Params Type LongVideoBench VideoMME
Val Long Medium Short Overall
Proprietary Models
GPT4-V - training-based 59.1 53.5 - - 59.9
GPT4-o - training-based 66.7 65.3 - - 71.9
Gemini-1.5-Pro - training-based 64.0 67.4 - - 75.0
Open-Sourse MLLMs
VideoLLaVA(Lin et al. 2023) 7B training-based 58.2 36.2 38.0 45.3 39.9
VITA-1.5(Fu et al. 2025) 7B training-based - 47.1 54.2 67.0 56.1
mPLUG-Owl3(Ye et al. 2024) 7B training-based 52.1 50.1 57.7 70.0 59.3
VideoLLaMA3(Zhang et al. 2025) 7B training-based 59.8 54.9 63.7 80.1 66.2
Oryx-1.5(Liu et al. 2024c) 7B training-based 56.3 - - - 58.3
QuoTA(Luo et al. 2025) 7B training-free 59.0 55.7 64.9 77.1 65.9
Video-XL(Shu et al. 2024) 7B training-based 49.5 - - - 55.5
ViLaMP(Cheng et al. 2025) 7B training-based 57.8 - - - 67.5
AdaReTaKe(Wang et al. 2025a) 7B traning-free 62.6 58.3 - - 67.7
AKS(Tang et al. 2025a) 7B training-free 62.7 - - - 65.3
Kangaroo(Liu et al. 2024a) 8B traning-based 54.8 - - - 46.6
NVILA(Liu et al. 2024d) 8B training-based 57.7 54.8 62.2 75.7 64.2
ByteVideoLLM(Wang et al. 2024a) 14B training-based - 56.4 62.9 74.4 64.6
Qwen2-VL(Wang et al. 2024b) 7B training-based 55.6 53.8 - - 63.3
Qwen2-VL w/ APVR(ours) 7B training-free 60.9 55.1 66.2 73.9 65.2
Qwen2.5-VL(Bai et al. 2025b) 7B training-based 59.5 55.6 - - 65.4
Qwen2.5-VL w/ APVR(ours) 7B training-free 64.9 59.1 67.8 78.3 68.4
Table 1: Video understanding accuracy(%) on LongVideoBench and VideoMME. The proposed APVR is applied upon two baseline MLLMs. The Methods are divided into two categories: training-based and training-free. APVR based on Qwen2.5-VL achieved state-of-the-art results on both the two benchmarks.
@@
3.3 Pivot Token Retrieval
@@
First, the retrieved pivot frames $\mathcal{F}_{pivot}$ are processed by a Vision Encoder $\mathcal{G}(\cdot)$ and a Projector $\mathcal{P}(\cdot)$ to generate dense visual token representations: $\mathcal{T}_{vis}=\mathcal{G}(\mathcal{P}(\mathcal{F}_{pivot}))=\{\mathcal{T}_{t}\}_{t=1}^{K}$. Our objective is to retrieve the pivot token $\mathcal{T}_{pivot}$. To this end, we introduce two components: query-aware multi-layer attention scoring and adaptive token selection.
@@
$A_{cross}^{l}=\mathrm{softmax}(\frac{q_{text}^{l}\cdot(k_{vis}^{l})^{T}}{\sqrt{d}})$ (9)
, where $A_{cross}^{l}\in R^{h\times d\times d_{q}\times d_{v}}$, $q_{text}^{l}$ denotes the query states derived from $\{Q,Des,Sem\}$ and $k_{vis}^{l}$ denotes the key state derived from the visual tokens $\mathcal{T}_{vis}$. Then the token-level attention scores of each layer are aggregated at the query dimension by:
@@
This formulation captures how much attention each key token receives across all query positions, providing a comprehensive measure of its importance for the current layer’s computations. The superscript $l$ will be omitted for concision.
@@
$\rho_{w}=\mathrm{min}(1.0,\sqrt{\frac{|\{j:a_{j}>0.01\cdot\mathrm{max}(a)\}|}{L_{w}}})$ (12)
, where $L_{w}$ is the chunk length. The final selection ratio is determined by:
$\gamma_{w}=\rho_{w}\cdot\eta_{w}$ (13)
Head-wise Soft Voting Based on the ratio $\gamma_{w}$, we can select the pivot token in each chunk by:
$\mathcal{Z}_{w}=\mathrm{TopK}(a_{w}\odot\mathcal{T}_{w},\gamma_{w}L_{w})$ (14)
@@
$\mathcal{Z}_{w}=\mathrm{TopK}((\sum_{j=1}^{h}\mathrm{softmax}(a_{w,j}))\odot\mathcal{T}_{w},\gamma_{w}L_{w})$ (15)
, where the softmax function calculates the per-head scores and gives a soft voting via normalization and sum. Finally, we update the KV cache as the selected pivot token in each layer:
@@
4 Experiments
4.1 Experimental setup
@@
We evaluate the performance of APVR in two popular benchmarks that contains hour-level videos: LongVideoBench(Wu et al. 2024). A benchmark for long-context video understanding, consisting of 3,763 videos with 6,678 annotated multiple-choice questions across 17 categories of referred reasoning questions. VideoMME(Fu et al. 2024). A multi-modal evaluation benchmark focused on fine-grained video understanding tasks requiring precise visual-textual alignment and detailed semantic comprehension. It comprises 900 videos totaling 256 hours, with 2,700 manually labeled complex multiple-choice question-answer pairs across 30 subfields. To highlight the importance of the pivot information retrieval for visual understanding, we do not use video subtitles for the two benchmarks.
@@
We integrate APVR with two baseline MLLMs, namely, Qwen2-VL-7B (Wang et al. 2024b) and Qwen2.5-VL-7B (Bai et al. 2025b). The MLLMs process a prompt involving the question, retrieved pivot video frames, and the multi-choice answer options to generate responses. We first densely extracted the raw frames from video with $fps=2$. Then, the sampled frames are iteratively scored using two basic visual models: CLIP (Radford et al. 2021) with ViT-B-16 (Dosovitskiy et al. 2020) backbone, and Grounding-DINO (Liu et al. 2024b) with Swin-T (Liu et al. 2021) backbone, based on the expanded semantic information generated by LLM. All evaluations are conducted on 8 NVIDIA A800 GPUs with 80GB memory, using LMMs-Eval (Zhang et al. 2024a) framework. Under the above setup, our APVR enables processing of up to 1024 frames, compared to the baseline 7B model’s capacity of 256 frames under optimal MLLM configurations.
4.2 Comparison to the State-of-the-Art
Quantitative results
We first compare the accuracy of video question answering between APVR with recent training-based or training-free methods. All results are summarized in Table 1. Upon the Qwen2-VL baseline, APVR improves the overall accuracy by 9.5% and 3.0% in LongVideoBench and VideoMME, respectively. Upon a stronger baseline Qwen2.5-VL, our APVR also improves the overall accuracy by 8.3% in LongVideoBench and 4.6% in VideoMME, respectively. Notably, APVR based on Qwen2.5-Vl reaches 64.9% accuracy in LongVideoBench, which is higher than the powerful proprietary models such as GPT4-V and Gemini-1.5-Pro. The overall performance of APVR also surpasses other training-based and training-free methods with larger parameter scales such as 8B and 14B. Another notable result is that our APVR reaches highest accuracy on long and medium video samples of VideoMME, confirming its unique capability to preserve critical spatio-temporal information in extended sequences.

### v3

@@ Abstract
Current multimodal large language models (MLLMs) struggle with hour-level video understanding, facing significant challenges not only in modeling the substantial information volume of long videos but also in overcoming the memory wall and resource constraints during both training and inference. Although recent training-free approaches have alleviated resource demands by compressing visual features, their reliance on incomplete visual information limits the performance potential. To address these limitations, we propose Adaptive Pivot Visual information Retrieval (APVR), a training-free framework that hierarchically retrieves and retains sufficient and important visual information. It breakthroughs the memory wall limitation via two complementary components: Pivot Frame Retrieval employs query expansion and iterative spatio-semantic confidence scoring to identify relevant video frames, and Pivot Token Retrieval performs query-aware attention-driven token selection within up to 1024 pivot frames. This dual granularity approach enables the processing of hour-long videos while maintaining semantic fidelity. Experimental validations on three different baseline MLLMs demonstrate significant performance improvements up to 9.5%, 4.6% and 9.7% on LongVideoBench, VideoMME and MLVU, respectively. APVR achieves state-of-the-art results for both training-free and training-based approaches.
Introduction
The proliferation of long-form video content necessitates robust understanding capabilities that extend beyond the current limitations of multimodal large language models (MLLMs). While existing video-based models (lin2023videollava; zhang2025videollama; bai2025qwen2.5vl; team2025kwai) demonstrate proficiency on short video sequences, their scalability for hour-level video remains challenging. Specifically, they encounter not only the memory wall but also the inherent dilution of semantic information across extensive temporal spans.
The core challenge for long video understanding with Video MLLMs lies in efficiently modeling and extracting sparse but critical semantic information across both temporal and spatial dimensions. Existing research generally falls into three paradigms: (1) uniform frame sampling, (2) sparse key-frame retrieval, and (3) sufficient frames processing while requiring computational tricks.
@@ Introduction
The first paradigm, uniform sampling, understands the entire video by evenly selecting frames across the sequence (bai2025qwen2.5vl; zhang2025videollama). As illustrated in Fig. 1(a), this strategy samples frames that are not relevant and pivot, leading to wrong answer generated by MLLMs. The second line of work attempts to address inefficiency by retrieving only $N$ sparse, query-relevant key frames from long videos (ataallah2024goldfish; tang2025adaptive-aks; guo2025logic; ye2025re; cheng2025scaling-vilamp; luo2025quota). This key-frame selection greatly reduces the computational load; however, as shown in Fig. 1(b), it neglects the temporal-semantic relationships between events and often fails in scenarios that require comprehensive temporal reasoning or long-range logical understanding. To improve temporal coverage, a third paradigm retrieves a denser set of $kN$ pivot frames (Fig. 1(c)) to capture relevant information. Meanwhile, it quickly encounters the memory wall when fed to MLLMs. Recent solutions leverage training-time trick, such as multimodal sequence parallelism (fu2025vita1.5; liu2024nvila) or feature compression (shu2024videoxl; liu2025videoxlpro), to handle longer sequences. However, these methods require resource-intensive multi-stage re-training and are tightly coupled to specific MLLM architectures, limiting their adaptability in a rapidly evolving models.
@@
The inherent trade-off between temporal coverage and computational feasibility has hindered existing methods from achieving comprehensive video understanding. To address this core challenge, we propose Adaptive Pivot Visual information Retrieval (APVR) framework, a training-free approach that integrates spatio-semantic frame retrieval with token retrieval and compression, as shown in Fig. 1(d). This two-stage retrieval mechanism concentrates computational resources on the most relevant visual information for the input query, thus enabling efficient, scalable and adaptable understanding of hour-long videos without retraining or encountering memory limitations.
@@
Specifically, at the frame level, our Pivot Frame Retrieval (PFR) component expands the original query into four types of semantic information: objects, descriptions, relations, and semantics. This expansion enables more comprehensive frame scoring from spatial and semantic perspectives using complementary visual models: CLIP for semantic similarity and Grounding-DINO for object detection and spatial reasoning. The framework incorporates temporal diffusion mechanisms to maintain temporal coherence and employs adaptive resampling strategies to efficiently and iteratively refine frame scoring and final selection. At the token level, our Pivot Token Retrieval (PTR) component extends the concept of semantic importance to fine-grained visual representations. This component leverages query-aware multi-layer attention scoring to identify the most relevant visual tokens to select and the irrelevant visual tokens to compress within selected frames, employing dynamic chunk-wise token selection and head-wise soft voting mechanisms to maintain both computational efficiency and semantic accuracy.
The main contributions of this work are as follows:
@@
Video MLLMs  Recent years have witnessed significant advances in multimodal large language models (MLLMs) for visual perception and understanding. As a critical modality in visual data analysis, video has also been integrated into MLLM frameworks. Notable efforts include Qwen2.5-VL (bai2025qwen2), InternVideo2.5 (wang2025internvideo2), VideoLlama3 (zhang2025videollama), and VideoChat-Flash (li2024videochat), which explore video understanding through a unified architecture consisting of three key components: (a) a video encoder for spatio-temporal feature extraction, (b) a projector for aligning visual features with linguistic embeddings, and (c) a video-aware LLM for multimodal reasoning. Due to the inherent complexity of video content (e.g., multi-object interactions, long-range temporal dependencies), existing MLLMs often fail to extract semantically salient information for video-centric downstream tasks requiring precise spatiotemporal reasoning, such as video question answering and spatiotemporal video grounding. Some works(bai2025qwen2; wei2025videorope) exploit to aggregate timestamps in the RoPE(su2024roformer). However, these methods exhibit significant performance degradation when extrapolating to videos exceeding their pre-training duration. Moreover, current architectures also exhibit limitations of inefficient cross-modal alignment between visual information and linguistic queries, particularly in long-form videos exceeding 5 minutes (li2024videomamba).
@@
Long Video Understanding  Recently advances in long video understanding can be broadly categorized into training-based and training-free paradigms. In the training-based paradigm, MLLMs are directly trained using large-scale long video datasets. The whole training process is usually divided into several stages to stably improve different components of MLLM, as well as to improve the ability to model different types of input visual data (li2024videochat; zhang2025videollama). In order to process extra-long videos, current methods explore two different strategies: sequence parallelism (chen2024longvila; longva; shen2025longvita) or visual feature compression (song2024moviechat; shen2024longvu; liu2025videoxlpro). For instance, LongVILA (chen2024longvila) adopts a five-stage training pipeline with a novel multi-modal sequence parallelism to efficiently learn from long videos. Video-XL-Pro (liu2025videoxlpro) builds a learnable module to generate compact and comprehensive tokens for long video understanding. In the training-free paradigm, long video understanding task is organized as an agent-like process (liu2025videomind; wang2024videoagent), where MLLMs serve as reasoning engine within the system. These training-free methods mainly focus on retrieving the most important information in video frames (tang2025adaptive; park2024too; ye2025re; guo2025logic; cheng2025scaling-vilamp) or visual tokens (wang2025adaretake; luo2025quota). For example, AKS (tang2025adaptive) adopts a plug-and-play module for keyframe selection and information pre-filtering for video-based MLLMs. QuoTA (luo2025quota) adopts an ante-hoc training-free module for query-aware visual token assignment and important information retrieval. Despite these advancements, few methods simultaneously explore frame and token optimization, which is addressed in this work through adaptive pivot information retrieval.
Method
@@
Overview
The framework of the proposed APVR is illustrated in Fig. 2, which is a training-free system. Given an input video $\mathcal{F}=\{f_{t}\}_{t=1}^{N}$ with $N$ frames sampled under specific fpsfps and query $Q$, the system aims to retrieve the most important visual information for accurate response. To achieve this, we build two main components in APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTR). In PFR, we search pivot frames $\mathcal{F}_{pivot}$ with high confidence scores using expanded queries and basic visual models. In PTR, we explore an attention-driven token selection in MLLMs to retrieve pivot tokens $\mathcal{T}_{pivot}$ that are most relevant with expanded queries.
Pivot Frame Retrieval
@@
Objects: The key objects $Obj_{k}$ or cue objects $Obj_{c}$ that are detectable by visual models such as Grounding-DINO (e.g., “bold person”,“phone”,“blue background”). All objects are aggregated as $Obj=\{Obj_{k},Obj_{c}\}$.
@@
[…]
@@
$\mathcal{S}_{i}=\mathrm{max}(\mathcal{S}_{i},\frac{\mathcal{S}_{t}}{1+|i-t|}),i\in[t-w,t+w].$ (2)
@@
$\mathcal{H}_{s}=\mathrm{TopK}(\mathcal{S}\odot\mathcal{U},\frac{N}{2\nabla_{p}}).$ (3)
@@
$\mathcal{E}_{\gamma}(s_{i})=-\sum_{j=i-\gamma}^{i+\gamma}e_{j}\cdot\text{log}(e_{j}),e_{j}=\frac{s_{j}}{\sum_{k=j-\gamma}^{j+\gamma}s_{k}},$ (4)
then the frames with high entropy are identified by:
$\mathcal{H}_{e}=\{f_{t}\mid\mathcal{E}_{\gamma}(s_{t})>\mu+0.5\sigma\},$ (5)
where $\mu$ and $\sigma$ are the mean and standard deviation of the entropy in $\mathcal{C}$.
@@
$\mathcal{C}_{\text{rand}}\sim\mathrm{Unif}(\mathrm{Perm}(S\odot\mathcal{C},\alpha N_{\mathcal{C}})),$ (8)
where $\alpha$ balances the multinomial score-based importance sampling $\mathcal{C}_{\text{multi}}$ and the diverse random sampling $\mathcal{C}_{\text{rand}}$ to capture the full semantic content of the video and also to avoid falling into local optimal.
@@
Methods Params Training LVB VideoMME(w/o sub.) MLVU
-free val Long Overall dev
Proprietary Methods
GPT4-V(openai2023gpt4v) - No 59.1 53.5 59.9 -
GPT4-o(hurst2024gpt4o) - No 66.7 65.3 71.9 64.6
Gemini-1.5-Pro(team2024gemini) - No 64.0 67.4 75.0 -
Open-Sourse Methods
VideoLLaVA(lin2023videollava) 7B No 58.2 36.2 39.9 67.0
VITA-1.5(fu2025vita1.5) 7B No - 47.1 56.1 -
mPLUG-Owl3(ye2024mplug) 7B No 52.1 50.1 59.3 63.7
Oryx-1.5(liu2024oryx) 7B No 56.3 - 58.3 67.5
QuoTA(luo2025quota) 7B Yes 59.0 55.7 65.9 71.9
Video-XL(shu2024videoxl) 7B No 49.5 - 55.5 64.9
ViLaMP(cheng2025scaling-vilamp) 7B No 57.8 - 67.5 72.6
AdaReTaKe(wang2025adaretake) 7B Yes 62.6 58.3 67.7 75.0
AKS(tang2025adaptive-aks) 7B Yes 62.7 - 65.3 -
NVILA(liu2024nvila) 8B No 57.7 54.8 64.2 70.1
ByteVideoLLM(wang2024dynamic-bytevideollm) 14B No - 56.4 64.6 65.0
Qwen2-VL(wang2024qwen2) 7B No 55.6 53.8 63.3 66.9
Qwen2-VL w/ APVR(ours) 7B Yes 60.9(+9.5%) 55.1(+2.4%) 65.2(+3.0%) 73.4(+9.7%)
Qwen2.5-VL(bai2025qwen2.5vl) 7B No 59.5 55.6 65.4 70.2
Qwen2.5-VL w/ APVR(ours) 7B Yes 64.9(+9.1%) 59.1(+6.3%) 68.4(+4.6%) 76.1(+8.4%)
VideoLLaMA3(zhang2025videollama) 7B No 59.8 54.9 66.2 73.0
VideoLLaMA3 w/ APVR(ours) 7B Yes 63.5(+6.2%) 58.7(+6.9%) 68.1(+2.9%) 77.2(+5.5%)
Table 1: Video understanding accuracy(%) on LongVideoBench (LVB), VideoMME and MLVU. We categorize the methods based on whether they are training-free. The propsed APVR achieved state-of-the-art results on all the three benchmarks and improves the performance based on three different baseline models. Bold denotes the best result while underline the second best. The relative improvement compared to the baseline models are also calculated and listed in the parenthesis.
@@
Pivot Token Retrieval
@@
First, the retrieved pivot frames $\mathcal{F}_{pivot}$ are processed by a Vision Encoder $\mathcal{G}(\cdot)$ and a Projector $\mathcal{P}(\cdot)$ to generate dense visual token representations: $\mathcal{T}_{vis}=\mathcal{P}(\mathcal{G}(\mathcal{F}_{pivot}))=\{\mathcal{T}_{t}\}_{t=1}^{K}$. Our objective is to retrieve the pivot token $\mathcal{T}_{pivot}$. To this end, we introduce two components: query-aware multi-layer attention scoring and adaptive token selection.
@@
$A_{cross}^{l}=\mathrm{softmax}(\frac{q_{text}^{l}\cdot(k_{vis}^{l})^{T}}{\sqrt{d}}),$ (9)
where $A_{cross}^{l}\in R^{h\times d\times d_{q}\times d_{v}}$, $q_{text}^{l}$ denotes the query states derived from $\{Q,Des,Sem\}$ and $k_{vis}^{l}$ denotes the key state derived from the visual tokens $\mathcal{T}_{vis}$. Then the token-level attention scores of each layer are aggregated at the query dimension by:
@@
This formulation captures how much attention each key token receives across all query positions, providing a comprehensive measure of its importance for the current layer. The superscript $l$ will be omitted for concision.
@@
$\rho_{w}=\mathrm{min}(1.0,\sqrt{\frac{|\{j:a_{j}>0.01\cdot\mathrm{max}(a)\}|}{L_{w}}}),$ (12)
where $L_{w}$ is the chunk length. The final selection ratio is determined by:
$\gamma_{w}=\rho_{w}\cdot\eta_{w}.$ (13)
Head-wise Soft Voting: Based on the ratio $\gamma_{w}$, we can select the pivot token in each chunk by:
$\mathcal{Z}_{w}=\mathrm{TopK}(a_{w}\odot\mathcal{T}_{w},\gamma_{w}L_{w}).$ (14)
@@
$\mathcal{Z}_{w}=\mathrm{TopK}((\sum_{j=1}^{h}\mathrm{softmax}(a_{w,j}))\odot\mathcal{T}_{w},\gamma_{w}L_{w}),$ (15)
where the softmax function calculates the per-head scores and gives a soft voting via normalization and sum. Finally, we update the KV cache as the selected pivot token in each layer:
@@
Method Long (10-60m) Avg
APVR (Qwen2-VL-7B) 51.2 60.9
w/o exp. sem. info. 50.4 59.1
APVR (Qwen2.5-VL-7B) 58.5 64.5
w/o exp. sem. info. 56.6 63.2
APVR (VideoLLaMA3-7B) 56.9 63.5
w/o exp. sem. info. 55.5 62.3
Table 2: Ablations of expanded semantic information.
Experiments
Experimental Setup
@@
We evaluate the performance of APVR in three popular benchmarks that contains hour-level videos: LongVideoBench(wu2024longvideobench): A benchmark consisting of 17 categories of referred reasoning questions. VideoMME(fu2024videomme): A multi-modal evaluation benchmark focused on fine-grained video understanding tasks. MLVU(zhou2025mlvu): A large-scale long video benchmark with a large wide of 9 distinct tasks. To highlight the importance of the pivot information retrieval for visual understanding, we do not use subtitles for all benchmarks.
@@
We integrate APVR with three baseline MLLMs, namely, Qwen2-VL-7B (wang2024qwen2) , Qwen2.5-VL-7B (bai2025qwen2.5vl) and VideoLLaMA3 (zhang2025videollama). The MLLMs process a prompt involving the question, retrieved pivot video frames, and the multi-choice answer options to generate responses. We first densely extracted the raw frames from video with $fps=2$. Then, the sampled frames are iteratively scored using two basic visual models with $\lambda=0.5$: CLIP (radford2021learning-clip) with ViT-B-16 (dosovitskiy2020image-vit) backbone, and Grounding-DINO (liu2024grounding-dino) with Swin-T (liu2021swinT) backbone, based on the expanded semantic information generated by LLM. All evaluations are conducted on 8 NVIDIA A800 GPUs with 80GB memory, using LMMs-Eval (zhang2024lmmseval) framework. Under the above setup, our APVR enables processing of up to $K=1024$ frames, compared to the baseline 7B model’s capacity of $256$ frames under optimal MLLM configurations. Our APVR can understand an hour-level long video and generate correct answers in up to 2 minutes.
Comparison to the State-of-the-Art
We first compare the accuracy of video question answering between APVR with recent training-based or training-free methods. All results are summarized in Table 1. Upon three different baseline MLLMs, APVR improves overall accuracy on all three widely used benchmarks. Notably, APVR based on Qwen2.5-VL reaches 64.9% accuracy in LongVideoBench, which is higher than the powerful proprietary models such as GPT4-V and Gemini-1.5-Pro. The overall performance of APVR also surpasses other training-based and training-free methods with larger parameter scales such as 8B and 14B. The complete verification on the benchmarks confirms the capability of APVR to preserve critical spatio-temporal information in long video sequences.